# Evaluating Evidence Attribution in Generated Fact Checking Explanations

### v2

@@ Abstract
Automated fact-checking systems often struggle with trustworthiness, as their generated explanations can include hallucinations. In this work, we explore evidence attribution for fact-checking explanation generation. We introduce a novel evaluation protocol, citation masking and recovery, to assess attribution quality in generated explanations. We implement our protocol using both human annotators and automatic annotators, and find that LLM annotation correlates with human annotation, suggesting that attribution assessment can be automated. Finally, our experiments reveal that: (1) the best-performing LLMs still generate explanations with inaccurate attributions; and (2) human-curated evidence is essential for generating better explanations. Code and data are available here: https://github.com/ruixing76/Transparent-FCExp.
@@
Rui Xing1   Timothy Baldwin1,2   Jey Han Lau1 1The University of Melbourne   2MBZUAI ruixing@student.unimelb.edu.au, tb@ldwin.net, jeyhan.lau@gmail.com
@@ 1 Introduction
Table 1: An example based on PolitiHop with claim $c$, veracity $v$, evidence passages $E$. $X$ is the generated explanation that cites $E$.
@@
A challenge in this task is that the explanations must provide accurate attribution of their key statements (Nakov et al., 2021). In journalistic fact-checking practice, journalists curate their sources and ensure the accuracy of every cited statement (Baker and Fairbank, 2022). Machine-generated explanations should this follow convention to be trustworthy. But even if generated explanations contain attribution, how do we assess whether the attribution is faithful to the source evidence? Current evaluation methods are primarily based on the Attribution to Identified Source (AIS) framework (Rashkin et al., 2023), whereby an important feature is “explicature”: the meaning of statement $s$ in linguistic context $c$. Generating explicatures typically involves either costly human annotation or a non-trivial process of decontextualization, meaning it is rarely used. Consequently, most automated evaluation methods directly use Natural Lanugage Inference (NLI) between $c$ and $s$ to quantify AIS (Gao et al., 2023; Liu et al., 2023; Li et al., 2024).
@@
To summarise, our contributions are: (1) we introduce a new evaluation protocol for evidence attribution in generated explanations via citation masking and recovery; (2) we automate our protocol via LLMs and compare it with NLI models, showing that LLMs align better with human annotators; and (3) in terms of evidence source, we explore the influence of human- and machine-selected evidence on explanation generation and find that human-selected evidence produces explanations that are more accurate in terms of evidence attribution.
@@ 3.1 Dataset
We use PolitiHop (Ostrowski et al., 2021) as our dataset for all experiments. Table 1 shows an example instance, comprised of a claim, a veracity label, a list of evidence passages that provide additional relevant information about the claim, and selected evidence – a subset of passages selected by human annotators that is most relevant to the claim and collectively clarifies the veracity of the claim. [^1] In PolitiHop, the evidence in each claim can sometimes have multiple subsets (e.g. {8,10} and {6, 7, 8}); for these cases we randomly select one subset for our experiments.
@@ 3.2 Automatic Explanation Generation
Given a claim, veracity label, and selected evidence (machine- or human-selected), we (zero-shot) prompt an LLM to generate an explanation to clarify the claim. The prompt explicitly asks it to use in-line citations, making it clear that all given evidence passages should be used (see “Generated Explanation” in Figure 1; the full prompt is given in Appendix Table 12). We test a range of LLMs in our experiments, noting that we always use the same LLM for evidence selection and explanation generation. We generate explanations for 100 claims sampled from PolitiHop.
@@
For the evidence selector and explanation generator, we experiment with three language models: GPT-4 (OpenAI, 2024), GPT-3.5, and LLaMA2 (Touvron et al., 2023). [^2] Detailed model version: GPT-4 = gpt-4-0613, GPT-3.5 = gpt-3.5-turbo, and LLaMA2 = Llama-2-7b-chat. Note that we always use the same LLM for both evidence selection and explanation generation, and as such the machine-selected evidence can also be interpreted as self-selected evidence. Table 2 presents data statistics for the explanation generation process. In general, we can see that the machine-selected evidence set tends to be larger (i.e. it includes more passages), and the explanations generated using machine-selected evidence are also longer.
@@ 3.3 Human Assessment Protocol
Assume we have a claim $c$, a veracity label $v$, $m$ evidence passages $E=\{e_{1},e_{2},e_{3},...,e_{m}\}$, and a generated explanation with $n$ sentences $X=\{x_{1},x_{2},x_{3},...,x_{n}\}$. [^3] Sentences are segmented with spaCy v3.7.2.
A subset of sentences $X_{cit}=\{x_{i},x_{i+1},...,x_{j}\}\subseteq{X}$ contains inline citations (e.g. [6]), [^4] Technically there is a possibility that $X=\emptyset$ when the LLM fails to cite any evidence passages, but we did not see this in practice. where ${1}\leq{i}\leq{j}\leq{n}$. We randomly select $e_{k}\in E$ (${1}\leq{k}\leq{m}$) and mask its inline citation marker in the explanation (e.g. you are wrong [6] $\rightarrow$ you are wrong), producing $X^{mask}$, the masked explanation. We denote $X^{mask}_{cit}\subseteq{X_{cit}}$ as the subset of explanation sentences where their citation markers are removed.
@@ 3.4 Human Annotation Details
Figure 3: The inner-annotator agreement of human annotation. “Evi Src” indicates the whether the evidence is human or machine selected.
@@
Each claim was annotated by 5 annotators. As agreement metrics like kappa (McHugh, 2012) are not applicable to our multi-label scenario, we use Krippendorff’s alpha (Krippendorff, 2011) to compute annotator agreement. [^7] For Krippendorff’s alpha, we use Jaccard distance to measure distance between two annotations (recall that each annotation result is a set, and as such we use Jaccard distance to measure the degree of set overlap).
Figure 3 presents the inter-annotator agreement results. On average, GPT-4 explanations produce the best agreement, followed by GPT-3.5. LLaMA2 has the lowest agreement, because its generated explanations tend to be quite noisy (we will revisit this in the next section). Interestingly, explanations generated using human-selected evidence have a lower agreement score. Overall, these agreement results indicate moderate to good agreement for GPT-4 and GPT-3.5.
@@ 3.5 Human Evaluation Results
Given the human annotation, we now compute citation recovery performance (set precision, recall, and F1 in Section 3.3) to assess attribution quality in the generated explanations. Results are presented in Table 3.
GPT-4 has the best performance by a comfortable margin, suggesting that its generated explanations cite their sources more accurately than the other LLMs. Manual analysis reveals that LLaMA2 explanations often contain incorrect citations, such as [1][2][3] which are directly copied from the prompt/instruction, as well as random links. That said, even the best model (GPT-4 with machine-selected evidence) only produces an F1 of approximately 0.74, which means a good proportion of citations are still not faithful. There’s also substantial variance ($\pm$0.31), indicating the performance is far worse in the worst case scenario.
@@ 4 Automatic Evaluation of Evidence Attribution
In terms of pretrained NLI models, we use two models: DeBERTaV3 (He et al., 2023) and TRUE (Honovich et al., 2022). [^8] Detailed model versions: DeBERTaV3 = deberta-v3-base, TRUE = t5_xxl_true_nli_mixture. The output label set for DeBERTaV3 is “entailment”, “neutral” and “contradiction”, and for TRUE is “1” (entailment) and “0” (not entailment). For our experiments we are only concerned with the entailment class.
@@ 4.1 Agreement with Human Annotation
To understand how reliable LLMs and NLI models are as automatic annotators, we first assess their agreement with human annotations. We compute the agreement between automated annotator and human annotator using krippendorff’s alpha in a similar manner to that described in Section 3.5. But here we compute the agreement between two sets of annotation results: (1) an automatic annotator’s results (e.g. GPT-4’s); and (2) aggregate results from all human annotators. [^10] That is, we take the union of all recovered sentences ($X_{pred}^{mask}$) for each human annotator into a larger set. We also standardise any labels outside the ground truth set ($X_{cit}^{mask}$) to a common label, as we are mostly interested in understanding the extent to which the automatic annotator selects the ground truth label when human annotators do so. [^11] For example, if the ground truth set ($X_{cit}^{mask}$) is {9}, and the predicted set ($X_{pred}^{mask}$) from the automatic and human (grouped) annotator is {7, 9} and {5, 9} respectively, we will post-process both predicted sets into {-2, 9}.
@@ 4.2 Automatic Evaluation Results
In Table 5, we see that most automatic annotators have high agreement with human annotators: GPT-4 has the best attribution quality, followed by GPT-3.5 and LLaMA2. That said, if we look at the magnitude of the attribution quality, GPT-3.5 seems to overestimate this (i.e. its F1 numbers are much higher than those for humans), and this correlates with the agreement performance we saw in Section 4.1 where GPT-3.5 has a lower agreement. GPT-4 produces results that best align with human annnotators, and this is also reflected in Section 4.1 where GPT-4 has the highest agreement.
@@
Looking at both Table 5 and Table 6, we see a similar trend across both sample and full, where GPT-4 is consistently the best generator, followed by GPT-3.5 and LLaMA2. Interestingly though, if we look at the impact of evidence source (“human” vs. “machine”), for sample we see a small improvement from “machine” to “human”, but this gap widens further in full (for generator GPT-3.5 and LLaMA2 in particular), suggesting that human-selected evidence does produce better explanations. This observation contrasts what we saw previously in Section 3.5 (where machine-selected evidence is on-par with human-selected evidence), and shows that it is important to assess attribution for every sentence in the explanation to get the complete picture.
To provide a more intuitive understanding of the overall transparency of the generated explanations, we present Table 7 to show the proportion of explanations where every sentence attributes the source accurately (evaluated by GPT-4). We observe a large gap between “human” and “machine”, again reinforcing that human-curated evidence leads to more transparent explanations.
Discussion
Putting these results together, we can conclude as follows: (1) GPT-4 is the best model for generating fact-checking explanations; (2) there is still much room for improvement for explanation generation, because even with human-curated evidence, only 31% of GPT-4’s explanations attribute the sources accurately; (3) GPT-4 is the best automatic annotator for assessing evidence attribution; and (4) the selection of evidence is crucial, and we found that human-selected evidence tends to produce higher quality explanations across different LLM generators.
@@ Discussion
Table 7: The proportion of Explanation where all sentences are transparent. We use a threshold of 0.6 for attribution F1 score to classify instance as “transparent”.
@@ 5 Conclusion
This work only evaluates the evidence attribution of the generated explanations, i.e. how faithful they are based on the sources. We want to note that this evaluation does not assess factuality, which checks whether an explanation is factually correct, that said, there are already many studies on factuality, and our evaluation is orthogonal to these studies. Additionally, while we tried our best to craft the prompts, there remains the question of whether evidence attribution can be further improved with more prompt engineering. The PolitiHop dataset may contain out-of-date information, given it was developed a few years ago. As such, this may conflict with the LLM’s parametric knowledge for certain claims. We saw this in our analyses where LLMs ignore evidence passages and generate statements based on their intrinsic knowledge. Also, the claims and evidence passages in PolitiHop are generally short single sentences. In practice, claims and evidence passages are likely to be longer.
Ethics Statement
@@ Appendix D Experiment Details
We ran all offline models on 4 Nvidia A100 GPUs in a data parallel fashion. Explanation generation with GPT-4 and GPT-3.5 takes around 2 hours and LLaMA2 takes around 4 hours. The cost of generating explanations was \$49 (US Dollar) with GPT-4 and \$10 (US Dollar) with GPT-3.5. We present full evidence attribution results for sample setting in
@@ Appendix E Scientific Artifacts
We list the licenses of different artifacts used in this paper: PolitiHop (MIT)  [^14] https://github.com/copenlu/politihop, Huggingface Transformers (Apache License 2.0)  [^15] https://github.com/huggingface/transformers, and spaCy (MIT)  [^16] https://github.com/explosion/spaCy. Our source code and annotated data will be under MIT license.

### v3

@@ Abstract
Automated fact-checking systems often struggle with trustworthiness, as their generated explanations can include hallucinations. In this work, we explore evidence attribution for fact-checking explanation generation. We introduce a novel evaluation protocol – citation masking and recovery – to assess attribution quality in generated explanations. We implement our protocol using both human annotators and automatic annotators, and find that LLM annotation correlates with human annotation, suggesting that attribution assessment can be automated. Finally, our experiments reveal that: (1) the best-performing LLMs still generate explanations with inaccurate attributions; and (2) human-curated evidence is essential for generating better explanations. Code and data are available here: https://github.com/ruixing76/Transparent-FCExp.
@@
Rui Xing1  Timothy Baldwin1,2  Jey Han Lau1 1The University of Melbourne, 2MBZUAI ruixing@student.unimelb.edu.au, tb@ldwin.net, jeyhan.lau@gmail.com
@@ 1 Introduction
Table 1: An example in PolitiHop with claim $c$, veracity $v$, evidence passages $E$. $X$ is the generated explanation that cites $E$.
@@
A challenge in this task is that the explanations must provide accurate attribution of their key statements (Nakov et al., 2021). In journalistic fact-checking practice, journalists curate their sources and ensure the accuracy of every cited statement (Baker and Fairbank, 2022). Machine-generated explanations should this follow convention to be trustworthy. But even if generated explanations contain attribution, how do we assess whether the attribution is faithful to the source evidence? Current evaluation methods are primarily based on the Attribution to Identified Source (AIS) framework (Rashkin et al., 2023), whereby an important feature is “explicature”: the meaning of statement $s$ in linguistic context $c$. Generating explicatures typically involves either costly human annotation or a non-trivial process of decontextualization, meaning it is rarely used. Consequently, most automated evaluation methods directly use Natural Language Inference (NLI) between $c$ and $s$ to quantify AIS (Gao et al., 2023; Liu et al., 2023; Li et al., 2024).
@@
To summarize, our contributions are: (1) we introduce a new evaluation protocol for evidence attribution in generated explanations via citation masking and recovery; (2) we automate our protocol via LLMs and compare it with NLI models, showing that LLMs align better with human annotators; and (3) in terms of evidence source, we explore the influence of human- and machine-selected evidence on explanation generation and find that human-selected evidence produces explanations that are more accurate in terms of evidence attribution.
@@ 3.1 Dataset
We use PolitiHop (Ostrowski et al., 2021) as our dataset for all experiments. Table 1 shows an example instance, comprised of a claim, a veracity label, a list of evidence passages that provide additional relevant information about the claim, and selected evidence – a subset of passages selected by human annotators that is most relevant to the claim and collectively clarifies the veracity of the claim. [^1] In PolitiHop, the evidence in each claim can sometimes have multiple subsets (e.g. {8,10} and {6, 7, 8}); for these cases we randomly select one subset for our experiments. We sampled 100 instances from PolitiHop.
@@ 3.2 Automatic Explanation Generation
Given a claim, its veracity label, and selected evidence (machine- or human-selected), we use a zero-shot prompt with LLMs to generate an explanation to clarify the claim. The prompt explicitly asks it to use in-line citations, making it clear that all given evidence passages should be used (see “Generated Explanation” in Figure 1; the full prompt is given in Appendix Table 12). We test a range of LLMs in our experiments, noting that we always use the same LLM for evidence selection and explanation generation. We generate explanations for all 100 sampled claims from PolitiHop.
@@
For the evidence selector and explanation generator, we experiment with three language models: GPT-4 (OpenAI, 2024), GPT-3.5, and LLaMA2 (Touvron et al., 2023). [^2] Detailed model version: GPT-4 = gpt-4-0613, GPT-3.5 = gpt-3.5-turbo, and LLaMA2 = Llama-2-7b-chat. We also tested other models including LLaMA2-7B, FlanT5-xxl (Chung et al., 2024), Falcon-30B (Almazrouei et al., 2023) and MPT (Team, 2023) but excluded them because these models generated repetitive content with fabricated citations. Note that we always use the same LLM for both evidence selection and explanation generation, and as such the machine-selected evidence can also be interpreted as self-selected evidence. Table 2 presents a statistical breakdown of the explanation generation process. In general, we can see that the machine-selected evidence set tends to be larger (i.e. it includes more retrieved passages), and the explanations generated using machine-selected evidence are also longer.
@@ 3.3 Human Assessment Protocol
Assume we have a claim $c$, a veracity label $v$, $m$ evidence passages $E=\{e_{1},e_{2},e_{3},...,e_{m}\}$, and a generated explanation with $n$ sentences $X=\{x_{1},x_{2},x_{3},...,x_{n}\}$. [^3] Sentences are segmented with spaCy v3.7.2. A subset of sentences $X_{cit}=\{x_{i},x_{i+1},...,x_{j}\}\subseteq{X}$ contains inline citations (e.g. [6]), [^4] Technically there is a possibility that $X=\emptyset$ when the LLM fails to cite any evidence passages, but we did not see this in practice. where ${1}\leq{i}\leq{j}\leq{n}$. We randomly select $e_{k}\in E$ (${1}\leq{k}\leq{m}$) and mask its inline citation marker in the explanation (e.g. you are wrong [6] $\rightarrow$ you are wrong), producing $X^{mask}$, the masked explanation. We then denote $X^{mask}_{cit}\subseteq{X_{cit}}$ as the subset of explanation sentences where their inline citation markers are removed.
@@ 3.4 Human Annotation Details
Figure 3: The inter-annotator agreement for the human annotations. “Evi Src” indicates the whether the evidence is Human- or Machine-selected.
@@
Each claim was annotated by 5 annotators. As agreement metrics like kappa (McHugh, 2012) are not applicable to our multi-label scenario, we use Krippendorff’s alpha (Krippendorff, 2011) to compute annotator agreement. [^7] For Krippendorff’s alpha, we use Jaccard distance to measure distance between two annotations (recall that each annotation result is a set, and as such we use Jaccard distance to measure the degree of set overlap). Figure 3 presents the inter-annotator agreement results. On average, GPT-4 explanations produce the best agreement, followed by GPT-3.5. LLaMA2 has the lowest agreement, because its generated explanations tend to be quite noisy (we will revisit this in the next section). Interestingly, explanations generated using human-selected evidence have a lower agreement score. Overall, these agreement results indicate moderate to good agreement for GPT-4 and GPT-3.5.
@@ 3.5 Human Evaluation Results
Given the human annotation, we now compute citation recovery performance (set precision, recall, and F1 in Section 3.3) to assess attribution quality in the generated explanations. Results are presented in Table 3. GPT-4 has the best performance by a comfortable margin, suggesting that its generated explanations cite their sources more accurately than the other LLMs. Manual analysis reveals that LLaMA2 explanations often contain incorrect citations, such as [1][2][3] which are directly copied from the prompt/instruction, as well as random links. That said, even the best model (GPT-4 with machine-selected evidence) only produces an F1 of approximately 0.74, which means a good proportion of citations are still not faithful. There’s also substantial variance ($\pm$0.31), indicating the performance is far worse in the worst case scenario.
@@ 4 Automatic Evaluation of Evidence Attribution
In terms of pretrained NLI models, we use two models: DeBERTaV3 (He et al., 2023) and TRUE (Honovich et al., 2022). [^8] Detailed model versions: DeBERTaV3 = deberta-v3-base, TRUE = t5_xxl_true_nli_mixture. The output label set for DeBERTaV3 is “entailment”, “neutral”, and “contradiction”, and for TRUE is “1” (entailment) and “0” (not entailment). For our experiments we are only concerned with the entailment class.
@@ 4.1 Agreement with Human Annotation
To understand how reliable LLMs and NLI models are as automatic annotators, we first assess their agreement with human annotations. We compute the agreement between automated and human annotators using krippendorff’s alpha in a similar manner to that described in Section 3.5. Here, however, we compute the agreement between two sets of annotation results: (1) an automatic annotator’s results (e.g. GPT-4’s); and (2) aggregate results from all human annotators. [^10] That is, we take the union of all recovered sentences ($X_{pred}^{mask}$) for each human annotator into a larger set. We also standardise any labels outside the ground truth set ($X_{cit}^{mask}$) to a common label, as we are mostly interested in understanding the extent to which the automatic annotator selects the ground truth label when human annotators do so. [^11] For example, if the ground truth set ($X_{cit}^{mask}$) is {9}, and the predicted set ($X_{pred}^{mask}$) from the automatic and human (grouped) annotator is {7, 9} and {5, 9} respectively, we will post-process both predicted sets into {-2, 9}.
@@ 4.2 Automatic Evaluation Results
In Table 5, we can see that most automatic annotators have high agreement with human annotators: GPT-4 has the best attribution quality, followed by GPT-3.5 and LLaMA2. That said, if we look at the magnitude of the attribution quality, GPT-3.5 seems to overestimate this (i.e. its F1 numbers are much higher than those for humans), and this correlates with the agreement performance we saw in Section 4.1 where GPT-3.5 has a lower agreement. GPT-4 produces results that best align with human annotators, and this is also reflected in Section 4.1 where GPT-4 has the highest agreement among all models.
@@
Looking at both Table 5 and Table 6, we see a similar trend across both sample and full, where GPT-4 is consistently the best generator, followed by GPT-3.5 and LLaMA2. Interestingly though, if we look at the impact of evidence source (“human” vs. “machine”), for sample we see a small improvement from “machine” to “human”, but this gap widens further in full (for generator GPT-3.5 and LLaMA2 in particular), suggesting that human-selected evidence does produce better explanations. This observation contrasts with what we saw previously in Section 3.5 (where machine-selected evidence is on-par with human-selected evidence), and shows that it is important to assess attribution for every sentence in the explanation to get the complete picture.
To provide a more intuitive understanding of the overall evidence attribution quality of the generated explanations, we present Table 7 to show the proportion of explanations where every sentence attributes the source accurately. We observe a large gap between “human” and “machine” across all generators, as well as all automatic annotators. This again reinforces that human-curated evidence leads to more transparent explanations.
@@
GPT-3.5 Human 66% 50% 44%
Machine 35% 28% 16%
LLaMA2 Human 45% 41% 31%
Machine 24% 14% 6%
LLaMA3.1 Human 34% 26% 25%
Machine 13% 9% 5%
Gemma Human 24% 14% 9%
Machine 5% 2% 3%
Mistral Human 19% 13% 11%
Machine 8% 5% 1%
Table 7: The proportion of Explanation where all sentences are “Fully Attributed”. We use an attribution F1 threshold of 0.6 to classify an instance as “Fully Attributed”. The “Evi Src” indicates whether evidence is Human- or Machine-selected, “Ann” the automatic annotators and “Gen” the explanation generators.
Discussion
Putting these results together, we can conclude as follows: (1) GPT-4 is the best model for generating fact-checking explanations based on evidences; (2) there is still much room for improvement for explanation generation, because even with human-curated evidence, only 31% of GPT-4’s explanations attribute the sources accurately; (3) GPT-4 is the best automatic annotator for assessing evidence attribution; and (4) the selection of evidence is crucial, and we found that human-selected evidence tends to produce higher quality explanations across different LLM generators.
@@ 5 Conclusion
This work only evaluates the evidence attribution of generated explanations, i.e. how faithful they are based on the sources. It is important to note that this evaluation does not assess factuality, which examines whether an explanation is factually correct. That said, there are already many studies on factuality, and our evaluation is orthogonal to these efforts. Additionally, while we tried our best to optimize the prompts, there remains the question of whether evidence attribution can be further improved with more prompt engineering. The PolitiHop dataset may contain out-of-date information, given it was developed a few years ago. As such, this may conflict with the LLM’s parametric knowledge for certain claims. We saw this in our analyses where LLMs ignore evidence passages and generate statements based on their intrinsic knowledge. Also, the claims and evidence passages in PolitiHop are generally short single sentences. In practice, claims and evidence passages are likely to be longer.
Ethical Statement
@@
Acknowledgments
We thank our reviewers for their valuable reviews and feedback. We also thank our annotators for their contribution. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ Appendix D Experiment Details
We ran all offline models on 4 Nvidia A100 GPUs in a data parallel fashion. Explanation generation with GPT-4 and GPT-3.5 takes around 2 hours and LLaMA2 takes around 4 hours. The cost of generating explanations was \$49 (US Dollar) with GPT-4 and \$10 (US Dollar) with GPT-3.5. We present full evidence attribution results for sample setting in Table 8 and full setting in Table 11.
@@ Appendix E Scientific Artifacts
We list the licenses of different artifacts used in this paper: PolitiHop (MIT), [^14] https://github.com/copenlu/politihop Huggingface Transformers (Apache License 2.0). [^15] https://github.com/huggingface/transformers Our source code and annotated data are released under MIT license.